# On the Fundamental Limits of LLMs at Scale

### v1

@@ 2 What Makes LLMs Hallucinate?
Hallucination in large language models arises from four interconnected sources (Figure 2): (1) Fundamental limits from computability theory, uncomputability, and statistical learning that prove hallucination is mathematically inevitable Xu et al. (2024e); Banerjee et al. (2025); (2) data-induced hallucinations from incomplete coverage, noise, long-tail distributions, temporal decay, and conflicting information in training corpora Huang et al. (2025a); Wang et al. (2023); (3) evaluation misalignment where benchmarks reward confident fabrication over calibrated uncertainty Xu et al. (2024a); Kirichenko et al. (2025); and (4) creativity-factuality trade-offs where mechanisms enabling creative generation necessarily increase hallucination risk Nguyen et al. (2024); Peeperkorn et al. (2024). Together, these establish hallucination not as a transient engineering problem but as an intrinsic property of probabilistic language models.
@@ 2.1 Fundamental Limits
Hallucination in LLMs is not merely an engineering artifact of insufficient data or suboptimal training; rather, it reflects intrinsic computational and statistical limits that no (present) architecture, scale, or optimization can overcome Xu et al. (2024e); Banerjee et al. (2025). We formalize these intrinsic boundaries through three complementary lenses: computability theory, which shows that no enumerable class of models can correctly answer all computable queries Peng et al. (2024); Karpowicz (2025); uncomputability, which demonstrates that certain problems lie beyond the reach of any algorithm Melo et al. (2025); and statistical learnability, which reveals that even learnable functions require sample complexity that often exceeds practical limits Asher et al. (2023); Su (2025); Khakhar et al. (2023); Goldblum et al. (2023). These results together establish hallucination as an inevitable feature of any learning system operating over open-ended domains Huang et al. (2025a); Dziri et al. (2023).
@@
We begin with the most fundamental limit: the impossibility of perfect learning for any enumerable collection of models. Modern LLMs, viewed as computable functions mapping input strings to output distributions, belong to computably enumerable sets (e.g., all polynomial-time Turing machines). A classical diagonalization argument originating from Cantor’s proof of uncountable infinities and adapted to learning theory reveals that any such enumerable collection must hallucinate on some inputs Tong et al. (2024b). We formalize this as follows.
@@
Beyond enumeration-based limitations, certain problems are undecidable: no algorithm can solve them for all inputs, regardless of computational resources. The canonical example is the Halting Problem, which asks whether a given computer program will finish running or continue to run forever on a given input. This was proven undecidable by Turing in 1936 Turing and others (1936). When an LLM encounters such queries, it faces an impossible dilemma: refusing to answer reveals incompleteness, while attempting an answer inevitably leads to hallucination Xu et al. (2024e). We formalize this inherent limitation.
@@
By assumption, $M^{\prime}$ correctly decides $f_{\mathrm{halt}}(\pi)$ for all $\pi\in\Pi$. Since $h$ is computable and the table lookup is computable, $M^{\prime}$ is a computable decider for the Halting Problem. This contradicts the undecidability of the Halting Problem Turing and others (1936). Therefore, $S_{h}$ must be infinite. ∎
Theorem 3 shows that undecidability creates an insurmountable barrier: no matter how sophisticated an LLM becomes, infinite failures are guaranteed on such problems. Moreover, variants of undecidable problems permeate practical applications: code analysis tasks (“Will this loop terminate?”), logical consistency checking (“Does this axiom set entail a contradiction?”), and self-referential queries (“Generate a sentence you cannot generate”) all inherit this fundamental impossibility Xu et al. (2024e). These are not contrived examples but questions users naturally pose to LLMs, making uncomputability-induced hallucination practically relevant. Yet even when we restrict attention to computable and decidable problems, statistical barriers remain.
@@
where $\widehat{\mathcal{R}}_{\text{hal}}(h)$ is the empirical hallucination rate on $n$ training samples, $d$ is the VC dimension of the hypothesis class, and the bound holds with probability at least $1-\delta$ Sahoo et al. (2024). For high-capacity models (large $d$) or distributions with long tails (rare facts appearing in $\ll n$ examples), the generalization term remains large even with low training error.
@@
Theorem 4 can essentially be translated as: when knowledge lacks structure (e.g., birthdates of millions of individuals, precise numerical constants, arbitrary historical events), sample requirements scale linearly with the number of facts. Real-world corpora, while vast, are finite and contain each rare fact only sparsely, making hallucination on long-tail queries much more statistically likely Su (2025).
@@
Mitigation strategies, including retrieval-augmented generation (oracle access), continual learning (adaptive capacity expansion), and constraint-based decoding can reduce hallucination in specific domains but cannot eliminate it universally Béchard and Ayala (2024). The takeaway is that: hallucination is an intrinsic property of learning systems operating over unbounded, open-ended query spaces, and any deployment of LLMs must account for this irreducible uncertainty.
@@ 2.2 Data-Induced Hallucinations
No training corpus, regardless of size, can encode all knowledge. The set of facts about the world grows continuously, while training datasets represent finite snapshots. Even static domains suffer from coverage gaps; i.e., rare entities, niche topics, and low-resource languages receive sparse representation, forcing models to interpolate or guess when queried about underrepresented content Onoe et al. (2022); Mousavi et al. (2024); Cheng et al. (2024).
@@
which, for any finite corpus and unbounded knowledge domain, satisfies $\rho_{\text{cov}}\to 0$ as $|\mathcal{K}_{\text{world}}|\to\infty$. Queries targeting the knowledge gap force the model to extrapolate from seen patterns to unseen facts, a process prone to plausible-sounding fabrications Wang et al. (2023). Further, let $q\in\mathcal{K}_{\text{world}}\setminus\mathcal{K}_{\text{train}}$ be a query about missing knowledge. The model generates an answer by maximizing:
@@
Training corpora are typically scraped from the internet or aggregated from diverse sources and naturally contain substantial noise, factual errors, and deliberate misinformation Sahoo et al. (2024). Web text includes unverified claims, outdated information, satirical content misinterpreted as factual, and adversarial misinformation. When LLMs ingest such data, they learn both correct and incorrect associations with no inherent process to distinguish truth from falsehood during pretraining.
@@
where $d$ is model capacity and $n$ is dataset size. [^1] This bound on the expected risk does not explicitly depend on the confidence parameter $\delta$, unlike high-probability bounds. It reflects the floor on average error imposed by label noise. This lower bound implies that even with infinite data ($n\to\infty$), a non-zero noise rate $\eta>0$ imposes a floor on hallucination risk. Studies quantifying this effect show that common web scrapes contain 2-3% demonstrably false factual claims ($\eta\approx 0.02$-$0.03$), conspiracy theories appear with non-negligible frequency, and even curated datasets like Wikipedia exhibit temporal inconsistencies and edit wars that encode conflicting information Zhang et al. (2025d). The model’s objective of “next-token prediction” treats all training text equally, optimizing likelihood rather than veracity:
@@
without regard to whether $y_{i}$ is factually correct. Consequently, frequently repeated misinformation can dominate the learned distribution, leading models to confidently reproduce falsehoods Dufour et al. (2024).
@@
Real-world knowledge follows a heavily skewed distribution; a small number of entities and facts appear frequently, such as major historical figures or popular scientific concepts, while the vast majority occur rarely. This long-tail phenomenon creates a memorization challenge: models must store millions of low-frequency facts with minimal reinforcement Sahoo et al. (2024).
Empirical evidence shows that LLM factual accuracy degrades sharply for tail entities (Figure 3a). For instance, when asked about individuals with fewer than 10 Wikipedia page views per day, GPT-4’s factual precision drops below 40%, compared to >>90% for highly popular entities Kandpal et al. (2023). The model has insufficient exposure to reliably encode these facts and instead generates plausible-sounding but incorrect details by analogy to more common patterns. This aligns with what Theorem 4 states, which in this context implies that without sufficient training examples for each rare fact, generalization bounds guarantee high error rates.
The memorization challenge is further complicated by interference: the model’s finite capacity means that learning new facts can overwrite or distort previously learned information, a phenomenon known as catastrophic forgetting in continual learning Gekhman et al. (2024). As models grow and training data expands, managing this interference becomes increasingly difficult, with rare facts being the most vulnerable to degradation.
@@
Training data represents a temporal snapshot, but the world evolves continuously. Let $t_{\text{cutoff}}$ denote the training data cutoff time and $t_{\text{query}}$ the query time. Facts that were true at $t_{\text{cutoff}}$ may become outdated by $t_{\text{query}}$. For example, political leaders change, scientific consensus shifts, companies merge or dissolve, and technologies become obsolete Zhu et al. (2024a). Define the temporal staleness of a fact $f$ as:
@@
which increases with $\Delta t:=t_{\text{query}}-t_{\text{cutoff}}$. For time-sensitive domains, empirical studies show $\tau(f)$ can exceed 0.5 within months, i.e., over half of time-dependent facts become stale (Figure 3b) Lazaridou et al. (2021). Models trained on pre-2023 data confidently assert information valid only in that historical context, producing temporally induced hallucinations when queried about current events Zhu et al. (2024b).
This issue is particularly worse for rapidly evolving domains. A model queried about “the current treatment for disease X” may generate a response reflecting outdated clinical guidelines from its training distribution, presenting obsolete information as current fact Lazaridou et al. (2021). Unlike incompleteness (which can be addressed via knowledge expansion), temporal decay may require continuous model updates. This presents a difficult trade-off: avoiding the “catastrophic forgetting” of past knowledge is important, yet retaining outdated information leads to temporal hallucinations. This process is costly and technically challenging, and thus requires a careful balance between preserving and updating information.
Worse, models lack explicit temporal reasoning. They cannot reliably distinguish between timeless facts (e.g., mathematical theorems, $\tau(f)\approx 0$) and time-dependent facts (e.g., “the president of the United States”, $\tau(f)\gg 0$). When training data contains multiple temporal versions of a fact without clear timestamp annotations, the model may conflate them, producing anachronistic or contradictory outputs Wang et al. (2023).
@@
In training corpora, different sources disagree on contested facts, present conflicting interpretations of events, or reflect divergent cultural perspectives, and thus, inherent contradictions arise. When such conflicts are unresolved in training, the model learns a superposition of contradictory beliefs, surfacing whichever aligns with the prompt’s framing or the model’s implicit biases Huang et al. (2025a).
For example, politically contentious topics have multiple narratives in web text. An LLM trained on this mixture may generate different “facts” depending on subtle prompt variations, revealing that it has not learned a single coherent world model but rather a distribution over conflicting models Zhang et al. (2025d). Unlike simple factual errors, the model has learned both correct and incorrect information, and query-time stochasticity determines which surfaces.
Additionally, systematic biases in training data—gender stereotypes, racial prejudices, and the perspectives of the dominant cultures in the training data—become encoded in model parameters Gallegos et al. (2024). When generating content about underrepresented groups or non-Western contexts, models often fall back on stereotypes or fabricate details consistent with biased training patterns. This produces culturally skewed hallucinations that reflect and amplify societal biases present in the data.
@@
A subtle but critical issue arises from the discrepancy between training and deployment distributions. During training, models see gold-standard prefixes $x_{<t}$ from the corpus and predict the next tokens. At inference, they generate text autoregressively, feeding their own predictions $\hat{x}_{<t}$ back as input. Let $p_{\text{data}}(x_{<t})$ be the training prefix distribution and $p_{\text{model}}(x_{<t})$ be the distribution induced by autoregressive generation. This exposure bias means the model never experiences its own errors during training, leading to compounding mistakes at test time Wang and Sennrich (2020). The distributional shift can be quantified via the KL divergence:
@@
where $\Delta_{\text{shift}}(t-t_{0})$ quantifies the compounding effect. Over long generations, these errors accumulate—a phenomenon observed in tasks like multi-hop reasoning and long-form summarization, where factual accuracy degrades as output length increases Wang and Sennrich (2020). The model has no mechanism to detect or correct these drifts, as it was never trained on its own potentially erroneous outputs.
@@
A growing concern nowadays is benchmark contamination. Test datasets inadvertently appearing in training corpora inflate performance estimates and obscure true hallucination rates Xu et al. (2024a). When models have seen test examples during training, they can memorize answers rather than learn generalizable reasoning, leading to brittle performance that fails on out-of-distribution queries. This creates an illusion of reduced hallucination that evaporates under “true” novel queries.
Related issues include data leakage from private or copyrighted sources, where models reproduce verbatim or near-verbatim text from training data, and model-generated data in training corpora, where synthetic text produced by earlier LLMs contaminates datasets for subsequent models. This latter issue, termed “model collapse” in some literature, can amplify hallucination patterns; if an earlier model’s factual errors are treated as ground truth in a later model’s training data, those errors become entrenched and harder to correct Shumailov et al. (2023).

### v2

@@ 2 What Makes LLMs Hallucinate?
Hallucination in large language models arises from four interconnected sources (Figure 2): (1) Fundamental limits from computability theory, uncomputability, and statistical learning that prove hallucination is mathematically inevitable (Xu et al., 2024e; Banerjee et al., 2025); (2) data-induced hallucinations from incomplete coverage, noise, long-tail distributions, temporal decay, and conflicting information in training corpora (Huang et al., 2025a; Wang et al., 2023); (3) evaluation misalignment where benchmarks reward confident fabrication over calibrated uncertainty (Xu et al., 2024a; Kirichenko et al., 2025); and (4) creativity-factuality trade-offs where mechanisms enabling creative generation necessarily increase hallucination risk (Nguyen et al., 2024; Peeperkorn et al., 2024). Together, these establish hallucination not as a transient engineering problem but as an intrinsic property of probabilistic language models.
@@ 2.1 Fundamental Limits
Hallucination in LLMs is not merely an engineering artifact of insufficient data or suboptimal training; rather, it reflects intrinsic computational and statistical limits that no (present) architecture, scale, or optimization can overcome (Xu et al., 2024e; Banerjee et al., 2025). We formalize these intrinsic boundaries through three complementary lenses: computability theory, which shows that no enumerable class of models can correctly answer all computable queries (Peng et al., 2024; Karpowicz, 2025); uncomputability, which demonstrates that certain problems lie beyond the reach of any algorithm (Melo et al., 2025); and statistical learnability, which reveals that even learnable functions require sample complexity that often exceeds practical limits (Asher et al., 2023; Su, 2025; Khakhar et al., 2023; Goldblum et al., 2023). These results together establish hallucination as an inevitable feature of any learning system operating over open-ended domains (Huang et al., 2025a; Dziri et al., 2023).
@@
We begin with the most fundamental limit: the impossibility of perfect learning for any enumerable collection of models. Modern LLMs, viewed as computable functions mapping input strings to output distributions, belong to computably enumerable sets (e.g., all polynomial-time Turing machines). A classical diagonalization argument originating from Cantor’s proof of uncountable infinities and adapted to learning theory reveals that any such enumerable collection must hallucinate on some inputs (Tong et al., 2024b). We formalize this as follows.
@@
Beyond enumeration-based limitations, certain problems are undecidable: no algorithm can solve them for all inputs, regardless of computational resources. The canonical example is the Halting Problem, which asks whether a given computer program will finish running or continue to run forever on a given input. This was proven undecidable by Turing in 1936 (Turing and others, 1936). When an LLM encounters such queries, it faces an impossible dilemma: refusing to answer reveals incompleteness, while attempting an answer inevitably leads to hallucination (Xu et al., 2024e). We formalize this inherent limitation.
@@
By assumption, $M^{\prime}$ correctly decides $f_{\mathrm{halt}}(\pi)$ for all $\pi\in\Pi$. Since $h$ is computable and the table lookup is computable, $M^{\prime}$ is a computable decider for the Halting Problem. This contradicts the undecidability of the Halting Problem (Turing and others, 1936). Therefore, $S_{h}$ must be infinite. ∎
Theorem 3 shows that undecidability creates an insurmountable barrier: no matter how sophisticated an LLM becomes, infinite failures are guaranteed on such problems. Moreover, variants of undecidable problems permeate practical applications: code analysis tasks (“Will this loop terminate?”), logical consistency checking (“Does this axiom set entail a contradiction?”), and self-referential queries (“Generate a sentence you cannot generate”) all inherit this fundamental impossibility (Xu et al., 2024e). These are not contrived examples but questions users naturally pose to LLMs, making uncomputability-induced hallucination practically relevant. Yet even when we restrict attention to computable and decidable problems, statistical barriers remain.
@@
where $\widehat{\mathcal{R}}_{\text{hal}}(h)$ is the empirical hallucination rate on $n$ training samples, $d$ is the VC dimension of the hypothesis class, and the bound holds with probability at least $1-\delta$ (Sahoo et al., 2024). For high-capacity models (large $d$) or distributions with long tails (rare facts appearing in $\ll n$ examples), the generalization term remains large even with low training error.
@@
Theorem 4 can essentially be translated as: when knowledge lacks structure (e.g., birthdates of millions of individuals, precise numerical constants, arbitrary historical events), sample requirements scale linearly with the number of facts. Real-world corpora, while vast, are finite and contain each rare fact only sparsely, making hallucination on long-tail queries much more statistically likely (Su, 2025).
@@
Mitigation strategies, including retrieval-augmented generation (oracle access), continual learning (adaptive capacity expansion), and constraint-based decoding can reduce hallucination in specific domains but cannot eliminate it universally (Béchard and Ayala, 2024). The takeaway is that: hallucination is an intrinsic property of learning systems operating over unbounded, open-ended query spaces, and any deployment of LLMs must account for this irreducible uncertainty.
@@ 2.2 Data-Induced Hallucinations
No training corpus, regardless of size, can encode all knowledge. The set of facts about the world grows continuously, while training datasets represent finite snapshots. Even static domains suffer from coverage gaps; i.e., rare entities, niche topics, and low-resource languages receive sparse representation, forcing models to interpolate or guess when queried about underrepresented content (Onoe et al., 2022; Mousavi et al., 2024; Cheng et al., 2024).
@@
which, for any finite corpus and unbounded knowledge domain, satisfies $\rho_{\text{cov}}\to 0$ as $|\mathcal{K}_{\text{world}}|\to\infty$. Queries targeting the knowledge gap force the model to extrapolate from seen patterns to unseen facts, a process prone to plausible-sounding fabrications (Wang et al., 2023). Further, let $q\in\mathcal{K}_{\text{world}}\setminus\mathcal{K}_{\text{train}}$ be a query about missing knowledge. The model generates an answer by maximizing:
@@
Training corpora are typically scraped from the internet or aggregated from diverse sources and naturally contain substantial noise, factual errors, and deliberate misinformation (Sahoo et al., 2024). Web text includes unverified claims, outdated information, satirical content misinterpreted as factual, and adversarial misinformation. When LLMs ingest such data, they learn both correct and incorrect associations with no inherent process to distinguish truth from falsehood during pretraining.
@@
where $d$ is model capacity and $n$ is dataset size. [^1] This bound on the expected risk does not explicitly depend on the confidence parameter $\delta$, unlike high-probability bounds. It reflects the floor on average error imposed by label noise. This lower bound implies that even with infinite data ($n\to\infty$), a non-zero noise rate $\eta>0$ imposes a floor on hallucination risk. Studies quantifying this effect show that common web scrapes contain 2-3% demonstrably false factual claims ($\eta\approx 0.02$-$0.03$), conspiracy theories appear with non-negligible frequency, and even curated datasets like Wikipedia exhibit temporal inconsistencies and edit wars that encode conflicting information (Zhang et al., 2025d). The model’s objective of “next-token prediction” treats all training text equally, optimizing likelihood rather than veracity:
@@
without regard to whether $y_{i}$ is factually correct. Consequently, frequently repeated misinformation can dominate the learned distribution, leading models to confidently reproduce falsehoods (Dufour et al., 2024).
@@
Real-world knowledge follows a heavily skewed distribution; a small number of entities and facts appear frequently, such as major historical figures or popular scientific concepts, while the vast majority occur rarely. This long-tail phenomenon creates a memorization challenge: models must store millions of low-frequency facts with minimal reinforcement (Sahoo et al., 2024).
Empirical evidence shows that LLM factual accuracy degrades sharply for tail entities (Figure 3a). For instance, when asked about individuals with fewer than 10 Wikipedia page views per day, GPT-4’s factual precision drops below 40%, compared to >>90% for highly popular entities (Kandpal et al., 2023). The model has insufficient exposure to reliably encode these facts and instead generates plausible-sounding but incorrect details by analogy to more common patterns. This aligns with what Theorem 4 states, which in this context implies that without sufficient training examples for each rare fact, generalization bounds guarantee high error rates.
The memorization challenge is further complicated by interference: the model’s finite capacity means that learning new facts can overwrite or distort previously learned information, a phenomenon known as catastrophic forgetting in continual learning (Gekhman et al., 2024). As models grow and training data expands, managing this interference becomes increasingly difficult, with rare facts being the most vulnerable to degradation.
@@
Training data represents a temporal snapshot, but the world evolves continuously. Let $t_{\text{cutoff}}$ denote the training data cutoff time and $t_{\text{query}}$ the query time. Facts that were true at $t_{\text{cutoff}}$ may become outdated by $t_{\text{query}}$. For example, political leaders change, scientific consensus shifts, companies merge or dissolve, and technologies become obsolete (Zhu et al., 2024a). Define the temporal staleness of a fact $f$ as:
@@
which increases with $\Delta t:=t_{\text{query}}-t_{\text{cutoff}}$. For time-sensitive domains, empirical studies show $\tau(f)$ can exceed 0.5 within months, i.e., over half of time-dependent facts become stale (Figure 3b) (Lazaridou et al., 2021). Models trained on pre-2023 data confidently assert information valid only in that historical context, producing temporally induced hallucinations when queried about current events (Zhu et al., 2024b).
This issue is particularly worse for rapidly evolving domains. A model queried about “the current treatment for disease X” may generate a response reflecting outdated clinical guidelines from its training distribution, presenting obsolete information as current fact (Lazaridou et al., 2021). Unlike incompleteness (which can be addressed via knowledge expansion), temporal decay may require continuous model updates. This presents a difficult trade-off: avoiding the “catastrophic forgetting” of past knowledge is important, yet retaining outdated information leads to temporal hallucinations. This process is costly and technically challenging, and thus requires a careful balance between preserving and updating information.
Worse, models lack explicit temporal reasoning. They cannot reliably distinguish between timeless facts (e.g., mathematical theorems, $\tau(f)\approx 0$) and time-dependent facts (e.g., “the president of the United States”, $\tau(f)\gg 0$). When training data contains multiple temporal versions of a fact without clear timestamp annotations, the model may conflate them, producing anachronistic or contradictory outputs (Wang et al., 2023).
@@
In training corpora, different sources disagree on contested facts, present conflicting interpretations of events, or reflect divergent cultural perspectives, and thus, inherent contradictions arise. When such conflicts are unresolved in training, the model learns a superposition of contradictory beliefs, surfacing whichever aligns with the prompt’s framing or the model’s implicit biases (Huang et al., 2025a).
For example, politically contentious topics have multiple narratives in web text. An LLM trained on this mixture may generate different “facts” depending on subtle prompt variations, revealing that it has not learned a single coherent world model but rather a distribution over conflicting models (Zhang et al., 2025d). Unlike simple factual errors, the model has learned both correct and incorrect information, and query-time stochasticity determines which surfaces.
Additionally, systematic biases in training data—gender stereotypes, racial prejudices, and the perspectives of the dominant cultures in the training data—become encoded in model parameters (Gallegos et al., 2024). When generating content about underrepresented groups or non-Western contexts, models often fall back on stereotypes or fabricate details consistent with biased training patterns. This produces culturally skewed hallucinations that reflect and amplify societal biases present in the data.
@@
A subtle but critical issue arises from the discrepancy between training and deployment distributions. During training, models see gold-standard prefixes $x_{<t}$ from the corpus and predict the next tokens. At inference, they generate text autoregressively, feeding their own predictions $\hat{x}_{<t}$ back as input. Let $p_{\text{data}}(x_{<t})$ be the training prefix distribution and $p_{\text{model}}(x_{<t})$ be the distribution induced by autoregressive generation. This exposure bias means the model never experiences its own errors during training, leading to compounding mistakes at test time (Wang and Sennrich, 2020). The distributional shift can be quantified via the KL divergence:
@@
where $\Delta_{\text{shift}}(t-t_{0})$ quantifies the compounding effect. Over long generations, these errors accumulate—a phenomenon observed in tasks like multi-hop reasoning and long-form summarization, where factual accuracy degrades as output length increases (Wang and Sennrich, 2020). The model has no mechanism to detect or correct these drifts, as it was never trained on its own potentially erroneous outputs.
@@
A growing concern nowadays is benchmark contamination. Test datasets inadvertently appearing in training corpora inflate performance estimates and obscure true hallucination rates (Xu et al., 2024a). When models have seen test examples during training, they can memorize answers rather than learn generalizable reasoning, leading to brittle performance that fails on out-of-distribution queries. This creates an illusion of reduced hallucination that evaporates under “true” novel queries.
Related issues include data leakage from private or copyrighted sources, where models reproduce verbatim or near-verbatim text from training data, and model-generated data in training corpora, where synthetic text produced by earlier LLMs contaminates datasets for subsequent models. This latter issue, termed “model collapse” in some literature, can amplify hallucination patterns; if an earlier model’s factual errors are treated as ground truth in a later model’s training data, those errors become entrenched and harder to correct (Shumailov et al., 2023).